I am a Mom trying desperately to help my son with his homework. It has been along time since I have done algebra. I am hoping you can help me. Here is his question: The sum of two numbers is 437. Twice the first number plus 5 times the second number is 1393. Find the two numbers.
Answer by LilSkittleMd(119)

First set up 2 equations
x+y=437
2x+5y=1393
Solve the first equation for x=
x=437-y
Plug this equation into the second equation
2(437-y)+5y=1393
Distribute the 2 to the values in the ()
874-2y+5y=1393
Combine like terms
874+3y=1393
Subtract 874 from both sides of the equation
3y=519
Divide both sides of the equation by 3
y=173
Now plug in the value of y into one of the original equations
x+173=437
Subtract 173 from both sides
x=264
The two numbers are 264 and 173